# DreamView: Injecting View-specific Text Guidance into Text-to-3D Generation

### v1

@@ 3.1 DreamView-2D for Text-to-image Generation
Specifically, we denote the text embeddings of the overall text prompt and view-specific text prompts given by the CLIP text encoder [37] as $\mathrm{E}^{t}_{o}$ and $\mathrm{E}^{t}_{v}$, respectively, where the superscript ${}^{t}$ denotes ‘text embedding’. Similarly, their global representations, i.e., class tokens are also output by the text encoder of CLIP, denoted as $\mathrm{CLS}^{t}_{o}$ and $\mathrm{CLS}^{t}_{v}$. The image embedding is denoted as $\mathrm{E}^{i}$. In each U-Net block, we measure the similarity between the image and text embeddings via
@@
[…]
@@ 4.1 Implementation Details
[…]
@@ 4.3 Text-to-3D Generation
Discussion with image-to-3D generation. In addition to text-to-3D generation, another stream of methods synthesizes 3D assets via lifting 2D images. We compare ours with them in Figure 8, where the reference image is generated by SD-XL using the overall text prompt. According to the results, although SD-XL can generate the concepts of the penguin, scarf, and crossbody bag, the crossbody bag is still overlooked when the generated image is lifted to a 3D object via Zero-1-to-3 XL [8] and SyncDreamer [28]. More importantly, these methods accept a single image for generating 3D assets, which inherently contradicts the diversity of appearance across viewpoints, making such singe-image-to-3D methods difficult to achieve customizable 3D generation.
Figure 8: Comparisons between DreamView-3D and image-to-3D methods. The reference image for performing image-to-3D is generated by SD-XL with the overall text.
@@ 5 Limitations and Conclusions
Acknowledgement. This work was supported partially by the Guangdong NSF Project (No. 2023B1515040025, 2020B1515120085), NSFC (U21A20471, U1911401). This work was also supported by Alibaba Group through Alibaba Innovative Research Program.
@@
[…]
@@
In Figures S10, S11 and S12, we compared the 3D generation results of 7 different methods under 3 text prompts containing customized content. We start from azimuth=0 degrees and show a total of 8 views at 45-degree intervals. These results demonstrate that other methods either suffer from the 3D inconsistent problem, e.g., multi-face and multi-foot problems, or cannot accurately generate specified content based on the text. In contrast, our method can include both 3D consistency and viewpoint customization capabilities.
@@
S7.2 More Results with General Prompts
@@
S7.3 Combing DreamView-2D with SJC Guidance
@@
S7.4 Applying the Injection Module to Other Text-to-image Models
@@
S7.5 Other Ways to Use View-specific Text Guidance
@@
The output text prompts and the 3D generation results are shown in the last row of Figure S17. The text prompts given by GPT-4 can successfully generate the penguin with a scarf and a crossbody bag, suggesting that using LLMs to generate view-specific texts may become a potential way to reduce user burden.

### v2

@@ 3.1 DreamView-2D for Text-to-image Generation
Specifically, we denote the text embeddings of the overall text prompt and view-specific text prompts given by the CLIP text encoder [36] as $\mathrm{E}^{t}_{o}$ and $\mathrm{E}^{t}_{v}$, respectively, where the superscript t denotes ‘text embedding’. Similarly, their global representations, i.e., class tokens are also output by the text encoder of CLIP, denoted as $\mathrm{CLS}^{t}_{o}$ and $\mathrm{CLS}^{t}_{v}$. The image embedding is denoted as $\mathrm{E}^{i}$. In each U-Net block, we measure the similarity between the image and text embeddings via
@@
where the ‘GAP’ denotes the global average pooling operation converting the image feature map to a representation vector, and ‘cos’ indicates the cosine similarity operator. Applying this equation to $\mathrm{CLS}^{t}_{o}$ and $\mathrm{CLS}^{t}_{v}$, we obtain the image-text similarity of the overall text (Simo) and the view-specific text (Simv). The magnitude between these two similarity values determines which guidance will be injected into the current block. For example, a larger Simo suggests that the current $\mathrm{E}^{i}$ absorbs more overall guidance than view-specific guidance, and thus we inject the view-specific one. This process can be formulated as:
@@ 4.1 Implementation Details
DreamView-2D. We train the DreamView-2D by combining our 3D rendered dataset and the 2D LAION dataset [39] to ensure 3D consistency, customization, and visual quality. The model is initialized by the SD-v2.1 model [37] and is trained on 16 V100 GPUs with a total batch size of 2,048. The learning rate is set to 1e-4. For each 3D object, we randomly select four orthogonal image views and resize them to $256\times 256$ to train the model, and the corresponding camera positions are normalized into a sphere. Regarding the inference, we adopt the DDIM [48] sampler with 50 sampling steps and a classifier-free guidance (CFG) scale of 7.5 for generating four image views simultaneously. The margin is randomly sampled from -0.1 to 0.1 during training and fixed at -0.025 during inference. Moreover, we also evaluate the model with our validation set, where we measure the CLIP-text score and CLIP-image score between the generated images and their corresponding texts and ground truth images, as well as the inception score for quality judgment.
@@ 5 Limitations and Conclusions
Acknowledgement. This work was partially supported by the Guangdong NSF Project (No. 2023B1515040025, 2020B1515120085) and NSFC (U21A20471, U1911401), and the Guangdong Basic and Applied Basic Research Foundation (2023A1515012974). This work was also supported by Alibaba Group through the Alibaba Innovative Research Program. We also thank Kun-Yu Lin, Yi-Xing Peng, and Yu-Ming Tang for their helpful discussions.
@@
    Training: We use a combined dataset consisting of our rendered dataset and a subset of the 2D LAION dataset [39]. In each training iteration, we select 3D data with a 70% probability and 2D data with a 30% probability. For each 3D object, we randomly select 4 orthogonal image views, and their captions are suffixed with “3D asset” if the raw captions do not contain the word ‘3D’ to distinguish from 2D data. We use 16 V100 to train the model. The batch size is 256 (4 objects, each object has 4 image views, across 16 GPUs), and the gradient is accumulated from 8 batches, forming a total batch size of 2,048. We use the AdamW [29] optimizer, and the learning rate is set to 1e-4. The image size is $256\times 256$. We randomly drop the text condition with a 10% probability. The margin is randomly sampled from -0.1 to 0.1 for each training iteration. We use SD-v2.1 as the model initialization, and the base network architecture is mostly the same as those used in [34, 45, 54, 52].
@@
In Figures S9, S10 and S11, we compared the 3D generation results of 7 different methods under 3 text prompts containing customized content. We start from azimuth=0 degrees and show a total of 8 views at 45-degree intervals. These results demonstrate that other methods either suffer from the 3D inconsistent problem, e.g., multi-face and multi-foot problems, or cannot accurately generate specified content based on the text. In contrast, our method can include both 3D consistency and viewpoint customization capabilities.
@@
Figure S12: Comparisons between DreamView-3D and image-to-3D methods. The reference image for performing image-to-3D is generated by SD-XL with the overall text.
S7.2 Discussion with Image-to-3D Generation
In addition to text-to-3D generation, another stream of methods synthesizes 3D assets via lifting 2D images. We compare ours with them in Figure S12, where the reference image is generated by SD-XL using the overall text prompt. According to the results, although SD-XL can generate the concepts of the penguin, scarf, and crossbody bag, the crossbody bag is still overlooked when the generated image is lifted to a 3D object via Zero-1-to-3 XL [8] and SyncDreamer [28]. More importantly, these methods accept a single image for generating 3D assets, which inherently contradicts the diversity of appearance across viewpoints, making such singe-image-to-3D methods difficult to achieve customizable 3D generation.
@@
S7.3 More Results with General Prompts
@@
S7.4 Combing DreamView-2D with SJC Guidance
@@
S7.5 Applying the Injection Module to Other Text-to-image Models
@@
S7.6 Other Ways to Use View-specific Text Guidance
@@
Table S2: Comparisons on hand-written and GPT-given prompts.
How to obtain CLIP Score ($\uparrow$) Inception
View-specific text Overall Text View Text Score ($\uparrow$)
Hand-written 31.1 32.1 14.5
GPT3.5-given 31.4 31.6 14.8
The output text prompts and the 3D generation results are shown in the last row of Figure S17. The text prompts given by GPT-4 can successfully generate the penguin with a scarf and a crossbody bag, suggesting that using LLMs to generate view-specific texts may become a potential way to reduce user burden. Besides, Table S2 is the quantitative result of text-to-image generation on the validation set between hand-written and GPT3.5-given prompts, showing that the GPT-given prompts are quantitatively on par with hand-written prompts.
A quick verification of prompts. We here provide a quick verification of whether the used prompts can achieve satisfactory customization. In practice, users can refine their prompts according to the text-to-image generation results, as the quality of 3D generation is highly correlated with image generation. Since the image generation is much faster than the 3D generation (3 seconds vs. 55 minutes), the prompt refinement process will not take long.